OpenStudy (anonymous):

Help? :c For the given function, find the vertical and horizontal asymptote(s) (if there are any).

OpenStudy (anonymous):

\[f(x) =\frac{ ((2x^2)+1) }{ ((x^2)-4) }\]

OpenStudy (anonymous):

@whpalmer4 I'm pretty sure -2 and 2, but my answer choices confuse me with the third choice.. A) None B) x = 2, y = 2y = 0 C) x = 2, x = -2, y = 2 D) x = 2, y = 2, y = 1

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@Loser66 Peas help me..

OpenStudy (loser66):

I don't know :( like you , there is no right choice to me

OpenStudy (anonymous):

Awhh. :C Thanks anyways, hun.

OpenStudy (loser66):

how?

OpenStudy (anonymous):

Huh?

OpenStudy (loser66):

y=2 is not an asymptote

OpenStudy (anonymous):

Tell me about it..

OpenStudy (whpalmer4):

C is correct. \(y=2\) is a horizontal asymptote. \(x = \pm 2\) are vertical asymptotes. See the attached graph. The vertical lines, which are an artifact of the graphing software, indicate the vertical asymptotes, and the purple line shows the horizontal asymptote. The horizontal asymptote is obtained by dividing the highest power term of the numerator by the highest power term of the denominator:\[y = \frac{2x^2}{x^2} = 2\]

OpenStudy (anonymous):

Oh!!! Thank you sooooo much @whpalmer4 . You're my hero. :P God bless you!